Suppose that the demand curve is Qd = 120 - 6p and the supply curve is Qs = 18p. Calculate the effects of a specific tax of t=2 per unit on the equilibrium, government tax revenue,consumer surplus, producer surplus,welfare and deadweight loss?

1
Expert's answer
2021-11-04T13:43:32-0400

Original equilibrium:

"Qs=Qd"

"120-6P=18P"

"P=5"

"120-6(5)"

"Q=90"

Finding new equilibrium:

Per tax reduces supplier by "t(P-t)"

Supply curve Qs"=18(P-2)=18P-36"

Qs"=" Qd

"18P-36=120-6P"

"24P=156"

"P=6.5"

"Qs=Qd=120-6(6.5)"

"Q=81"

Prices that producers receive at new equilibrium "=" "6.5-2=4.5"

Government tax revenue=Amount traded"\u00d7" taxes

"=81\u00d72=162"

CS and PS is calculated from the area represented on the supply and demand graph

Consumer Surplus.

"\frac {1}{2}\times 81\times 13.5=546.75"

Producer Surplus

"\frac{1}{2}\times 4.5\times 81=182.25"

Welfare and deadweight loss is calculated from the triangle area between the sandwich of original demand and supply curve

"=\frac{1}{2}×(6.5-4.5)×(90-81)=9"